–4 times a number plus 12 is –116. Find the number
Answer by TimothyLamb(4379)

---
-4x + 12 = -116
-4x = -116 - 12
x = ( -116 - 12 ) / -4
x = 32
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php